# Should you use LLMs to simulate opinions? Quality checks for early-stage deliberation

### v1

@@ Abstract
The array of emergent capabilities of large language models (LLMs) has sparked interest in assessing their ability to simulate human opinions in a variety of contexts, potentially serving as surrogates for human subjects in opinion surveys. However, previous evaluations of this capability have depended heavily on costly, domain-specific human survey data, and mixed empirical results about LLM effectiveness create uncertainty for managers about whether investing in this technology is justified in early-stage research. To address these challenges, we introduce a series of quality checks to support early-stage deliberation about the viability of using LLMs for simulating human opinions. These checks emphasize logical constraints, model stability, and alignment with stakeholder expectations of model outputs, thereby reducing dependence on human-generated data in the initial stages of evaluation. We demonstrate the usefulness of the proposed quality control tests in the context of AI-assisted content moderation, an application that both advocates and critics of LLMs’ capabilities to simulate human opinion see as a desirable potential use case. None of the tested models passed all quality control checks, revealing several failure modes. We conclude by discussing implications of these failure modes and recommend how organizations can utilize our proposed tests for prompt engineering and in their risk management practices when considering the use of LLMs for opinion simulation. We make our crowdsourced dataset of claims with human and LLM annotations (TopicMisinfo) publicly available for future research.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant appeal and versatility across a wide range of tasks. While the primary training objective of the base model is to predict the next most likely token, training on vast and diverse datasets, combined with increasingly complex system architectures has resulted in the “emergence” of capabilities that were not explicitly anticipated during development [1]. For instance, LLMs have been shown to learn new tasks through in-context learning [2], perform coding tasks with remarkable accuracy [3], and have displayed human-like capabilities across different tasks [4, 5].
@@
However, existing empirical findings are mixed regarding the suitability of LLMs for human opinion simulation. Some studies indicate that LLMs prompted with specific demographic characteristics produce accurate predictions of group-level opinions [16, 19], while other studies show that LLMs prompted with demographic characteristics can fail to adequately represent group-level opinions, reporting particularly on inadequate performance for minority groups and non-Western identities [9, 20, 6, 21]. Findings have also shown that LLMs tend to flatten within-group differences of opinion [18, 22, 21], and may not generalize across different opinion tasks and domains [13, 23].
These mixed results indicate that, while there is a possibility that current or future iterations of LLMs could be leveraged to successfully simulate opinions—a particularly valuable prospect when surveys expose human subjects to harmful content [18] (e.g., content moderation) or trade secrets (e.g., marketing new products)—thorough evaluation is needed before relying on LLM outputs for any such application. However, current methodologies to evaluate LLM-based opinion simulation rely heavily on costly human data [16, 19, 9, 20, 18]. These costs present challenges to initial research and development efforts of LLMs for automating opinion surveys for two main reasons. First, given the existing uncertainty around whether LLMs effectively simulate opinions, there may be little incentive to commit financial resources up front. Second, paradoxically, if using simulated opinions in a pilot study requires human data to validate any conclusions, one may as well simply use human data. These challenges are exacerbated by the dynamic evolution of human opinion, as well as LLM-generated ones due to system updates.
To address the challenge of costly reliance of existing validation methodologies on human data, we develop a series of tests for early-stage deliberation that do not require human data for evaluation and instead probe LLM outputs for other desirable qualities. Specifically, we propose tests to evaluate whether LLMs demonstrate:
@@
    Model Stability Across Updates—whether models remain relatively consistent in their opinion outputs over time across minor updates;
@@
These tests are informed by research on assessing quality standards and biases in subjective data annotation [24, 25, 26, 27, 28, 29, 30, 31], and they are intended to help practitioners decide whether additional investment in human data is justified and guide the iterative design of prompts intended to meet these quality standards. We show that these proposed quality control tests, despite their simplicity, reveal crucial failure modes of state-of-the-art, commercially available LLMs, indicating that these are not yet reliable means to simulate group-level opinions. We conclude by discussing implications of the observed failure modes and recommend how organizations can utilize our proposed tests for prompt engineering and in their risk management practices when considering the use of LLMs for opinion simulation. We make our crowdsourced dataset of claims with human and LLM annotations (TopicMisifo) publicly available for future research.
Experimental Dataset
Content moderation is often cited as an appealing application for LLMs. This is because moderation tasks can be psychologically harmful for human workers [32], occur at high frequency [33], and demand the scalability that automated systems can provide [34]. We explore automating opinion surveys in the context of fact-checking workflows, where LLM-based opinion surveys could help scale how claims are prioritized for human verification. Similar to many other survey contexts, in this content moderation tasks the social background and demographic identities of human annotators can crucially shape their assessments [30, 35, 36]. It is critical that AI-driven systems should capture relevant variations in opinions across demographic groups, which can reflect a range of group-specific priorities and perspectives. Otherwise, it risks systematically overlooking the concerns of certain demographic groups, leaving them disproportionately vulnerable to misinformation harms [37]. Indeed, the risk of overlooking harms to certain demographic groups has already been observed among fact-checkers employing advanced claim prioritization tools and technologies [38, 39, 40].
To provide an empirical application of the proposed evaluation methodologies, we created the TopicMisinfo dataset, which consists of 160 claims—some true, some false—spanning a wide range of topics. These include potentially demographically divisive issues such as abortion and illegal immigration, as well as topics like entertainment and weather, which are generally associated with broader consensus. Anchored on results from a survey from the American National Election Studies (ANES) [41], we selected topics based on their likelihood to provoke disagreement or agreement about whether they warrant fact-checking, considering four key criteria: (i) perceived importance for prioritization, (ii) general public interest, (iii) potential harm to specific groups, and (iv) perceived factual validity. The focus of our study is on evaluating LLMs ability to simulate gender perspectives on claims on the third axis — potential harm of claims to specific groups — as research and past polling has pointed out that gender differences tend to be particularly salient for this item [42, 43].
We apply the proposed evaluation framework to a diverse set of commercially available LLMs: GPT-3.5-Turbo (checkpoints 06-13-23, 11-06-23, and 01-25-24), GPT-4, LLaMA 3, Titan-Text-Premier, and Mistral-Large—hereafter referred to as GPT3.5a, GPT3.5b, GPT3.5c, GPT-4, LLaMA3, Titan, and Mistral, respectively. To simulate group-level responses, we adopt the “backstory prompting” methodology [16], instructing models to respond from the perspectives of either men or women. These prompts build on established techniques for gender-specific language model conditioning [16, 20]. Full prompt details are available in Supplementary Materials S3. To assess alignment with stakeholder expectations—one of our three proposed evaluation criteria—we use two complementary strategies. First, we incorporate “gold standard” items inspired by comprehension-based quality control protocols in human annotation [27] to evaluate baseline output reliability for this task in Experiment 3a. Second, we incorporate crowdsourced annotations to demonstrate how human data—when available—can inform domain knowledge and serve as a valuable comparison point, which we address in Experiment 3b. See Supplementary Materials S1 and S2 for further information on dataset construction, examples of claims used in the opinion survey, and descriptive statistics of human annotations.
@@ 2 Results
In this section, we analyze the results of three quality checks for a particular content moderation task: predicting the harm of potential misinformation to certain groups as a means to prioritize it for review [^1] While not the only criteria relevant for prioritization, research has noted that this is an important axis of consideration used by fact-checkers [44].. These tests, designed to emphasize logical consistency, model stability, and alignment with stakeholder expectations, serve as sanity-checks in early-stage deliberation when considering the use of LLMs to simulate group-level opinions.
@@
Figure 1: Experiment 1 Results. Plausible ranges of $q_{0}$ across topics for different models are highlighted in blue.
Testing for logical consistency can be an effective way to gauge the reliability of human annotators [29] or data aggregation methods [45]. For instance, Arguete et al. [29] use logical contradiction as a measure of human annotator inattention. We introduce a related methodology for evaluating logical consistency of LLMs in group-level opinion simulation by testing whether a model, when prompted for an “average" or “neutral” perspective on a Likert-scale assessment, generates a stance that is a convex combination of its predicted group-level opinions. If a model’s average opinion response is positioned plausibly within the convex hull of divergent group opinions, it suggests a structured and internally consistent approach to generating viewpoints.
In this paper, we define statistical neutrality as an opinion located within the convex hull of all group-level opinions within a social category (e.g., gender, race, etc.), including the extrema (i.e., only representing one group). This is distinct from neutrality focused on disinterest or impartiality [46, 47]. However, there is no single established approach to prompting an LLM to generate a “statistically neutral” viewpoint. To address this, four distinct prompts were evaluated, each with different approaches to querying statistical neutrality. See Supplementary Materials S3 for exact prompt specifications.
@@
Experiment 2: Model Stability Over Time
Measuring the stability of a single LLM’s outputs across minor model updates (or “checkpoints”) offers critical insight into its reliability. This kind of test is motivated by quality control in human annotation pipelines, where inconsistent judgments cast doubt on an annotator’s credibility [28]. As Abercrombie et al. [25, pg.1] observe, “consistency and label stability are important because, without them, annotation schemes are unlikely to be repeatable or reproducible.”
Model stability is also closely tied to the concept of backwards compatibility, a long-standing concern in software engineering [48], and one of growing importance when we consider the use of “off-the-shelf” machine learning models applied to a variety of application domains [49]. In AI systems, backwards compatibility refers to whether updated models make categorically similar predictions or errors to previous versions. For LLMs, it is reasonable to expect backwards compatibility within a given “model class” (e.g., across checkpoints of GPT-3.5-Turbo), though not necessarily between model classes (e.g., GPT-3.5-Turbo vs. GPT-4). Updates within a model class are generally intended to improve safety and reliability, not to introduce significant behavioral or quality changes [50]. Yet, emerging evidence suggests this expectation is not always met. For example, Chen et al. [51] found that two checkpoints of GPT-3.5-Turbo disagreed on 27% of multiple-choice responses in the OpinionQA dataset.
@@
Figure 2: Experiment 2 Results. This chart highlights statistically significant changes in model outputs across three updates (June ‘23, Nov ‘23, and Jan ‘24) for three types of prompts: neutral, men-conditioned, and women-conditioned.
@@
Ratings of these quality control items should remain stable and low over time. These claims were all designed to be harmless truths or lies. However, we see significant differences in ratings across checkpoints for men-conditioned ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 0.19, p $<$ 0.001), neutral ($\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5c) = 0.55, p $<$ 0.001), and women-conditioned prompts ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 0.51, p $<$ 0.001).
@@
Table 1: Salient Examples of Opinion Differences Diverging from Priors in Experiment 3a. $\mu(c_{i})^{gender}_{LLM}$ represents the average predicted harmfulness rating for the claim for a given gender.
@@
Stakeholders that rely on LLM-simulated opinions for decision making have a vested interest in verifying that LLMs accurately reflect group-level viewpoints. If an LLM does not position group-level opinions appropriately, it risks introducing biases into human decision-making that depends on these model outputs. For example, if fact-checking systems rely on LLMs to estimate which claims different groups are likely to find harmful, a model that underrepresents gender differences in concern may deprioritize fact-checks on topics that disproportionately affect one group’s information environment. Notably, even in the absence of “ground truth" human data to validate a particular opinion survey task, stakeholders often have valuable domain knowledge that can be leveraged to validate the reliability of LLMs. In this experiment, we focus on the positioning of gender differences in opinion, asking: Do models position men’s and women’s opinions in ways that align with stakeholder expectations regarding the degree of disagreement across different topics?
We define “expectations” by leveraging topics where prior research, survey results, or common sense strongly indicates minimal or substantial disagreement between gender groups. Experiment 3 examines two complementary scenarios: (a) whether the model violates well-established priors from common sense, domain knowledge, or survey data—e.g., that a neutral statement like “A circle is round” should not elicit divergent opinions by gender, whereas a claim about abortion might, drawing inspiration from prior comprehension-based quality control checks [27]; and (b) whether the model distorts stakeholder expectations on less obviously gender-polarized issues (e.g., politics, immigration, health and science), by introducing task-specific human survey data that serves as a reasonable prior for stakeholder expectations. This two-part design allows us to identify when a model introduces artificial rifts or masks genuine ones, both crucial factors in gauging an LLM’s trustworthiness for representing group-specific viewpoints. See Materials and Methods for details on the methodology for Experiments 3a and 3b.
@@
To test for exaggeration of opinion differences, we compared the output of LLMs prompted to reflect the opinions of men and women on topics where we had very strong priors that no division should exist. For instance, the Gold topic in the TopicMisinfo dataset contains a set of items that are harmless and uncontroversially true (e.g., “A circle is round”) or false (e.g., “The tallest tree on Earth creates a land bridge to Mars”), and they were compiled and used as a means to make sure that human annotators were paying attention and comprehending the task. We would very reasonably expect no opinion disagreement on claims of this type. However, we see GPT3.5a ($\hat{L}_{GPT3.5a}=0.25$, p $<$ 0.001), GPT3.5b ($\hat{L}_{GPT3.5b}$ = 0.20, p $<$ 0.05), and GPT3.5c ($\hat{L}_{GPT3.5c}$ = 0.38, p $<$ 0.01), as well as Titan ($\hat{L}_{Titan}$ = 0.19, p $<$ 0.01), tended to produce statistically significant differences in opinions by gender for claims on this topic. See Table 1 for specific examples of exaggeration.
@@
When preliminary data from real human responses is available, this can inform the stakeholder prior. For this purpose, we use crowdsourced annotations of the TopicMisinfo dataset. When comparing LLM responses to crowdsourced human responses, GPT3.5a, GPT3.5b, and GPT3.5c tended to exaggerate differences between groups, particularly for topics that are not typically considered divisive between genders. For example, claims from topics such as entertainment ($\hat{L}_{GPT3.5a}=0.84$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.49$, $p<0.01$; $\hat{L}_{GPT3.5c}=0.55$, $p<0.001$), health & science ($\hat{L}_{GPT3.5a}=1.05$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.52$, $p<0.01$), U.S. politics ($\hat{L}_{GPT3.5a}=1.09$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.58$, $p<0.001$; $\hat{L}_{GPT3.5c}=0.36$, $p<0.01$), and weather & climate ($\hat{L}_{GPT3.5a}=0.65$, $p<0.01$; $\hat{L}_{GPT3.5b}=0.62$, $p<0.01$; $\hat{L}_{GPT3.5c}=0.46$, $p<0.05$) saw significant overestimation of group-level disagreement by these models.
@@
To test for erosion of opinion differences, we compared the output of LLMs prompted to reflect the opinions of men and women on topics where we had very strong priors that division should exist. For instance, it is well-established that men and women may perceive harms differently by claims about abortion and women’s health [43, 42], and this topic is included in the TopicMisinfo dataset. Thus, we have some prior that a reasonable LLM should display differences in outputs when conditioned to respond as if it were either a man or woman. However, we didn’t find any significant differences between the outputs from men- or women-conditioned prompts on the topic for Mistral ($\hat{L}_{Mistral}=0.26$, p = 0.09), GPT-4 ($\hat{L}_{GPT4}=0.06$, p = 0.16), and LLaMa3 ($\hat{L}_{LLaMa}=0.06$, p = 0.23). See Table 1 for specific examples of erosion.

### v2

@@ Abstract
The emergent capabilities of large language models (LLMs) have sparked interest in assessing their ability to simulate human opinions in a variety of contexts, potentially serving as surrogates for human subjects in opinion surveys. However, previous evaluations of this capability have depended heavily on costly, domain-specific human survey data, and mixed empirical results about LLM effectiveness create uncertainty for managers about whether investing in this technology is justified in early-stage research. To address these challenges, we introduce a series of quality checks to support early-stage deliberation about the viability of using LLMs for simulating human opinions. These checks, designed to evaluate Likert-scale opinion data, emphasize logical consistency, prompt stability, and alignment with stakeholder expectations of model outputs, thereby reducing dependence on human-generated data in the initial stages of evaluation. We demonstrate the usefulness of the proposed quality control tests in the context of AI-assisted content moderation, an application that both advocates and critics of LLMs’ capabilities to simulate human opinion see as a desirable potential use case. None of the tested models passed all quality control checks, revealing several failure modes. We conclude by discussing implications of these failure modes and recommend how organizations can utilize our proposed tests in their risk management practices when considering the use of LLMs for opinion simulation. We make our crowdsourced dataset of claims with human and LLM annotations (TopicMisinfo) publicly available for future research.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant appeal and versatility across a wide range of tasks. While the primary training objective of the base model is to predict the next most likely token, training on vast and diverse datasets, combined with increasingly complex system architectures, has resulted in the “emergence” of capabilities that were not explicitly anticipated during development [1]. For instance, LLMs have been shown to learn new tasks through in-context learning [2], perform coding tasks with remarkable accuracy [3], and have displayed human-like capabilities across different tasks [4, 5].
@@
However, existing evidence is inconclusive on the extent to which LLM-based methods can accurately simulate human opinions. Some studies show that including a demographic “backstory” in a prompt [15, 18, 19] (i.e., “backstory” prompting), fine-tuning [20], and in-context learning [21] are approaches that can successfully approximate group-level opinions, while others document systematic failures of these same approaches—especially when simulating the opinions of minority or non-Western groups [9, 22, 6, 23, 24]. Additional work finds that backstory prompting approaches—the most prevalent approach to simulating opinions in prior research—flattens within-group variance of opinions [17, 25, 23] and fails to generalize across topics [13, 26]. (See Supplementary Materials S0 for a comprehensive review.)
The mixed results in the empirical literature suggest that the reliability of LLM-simulated opinions cannot be assumed across domains and instead requires rigorous validation for each new use case. Importantly, however, current validation methods rely on gathering high-quality human-labeled data—the cost of which can rival that of traditional surveys. For early-stage projects, this upfront expense is likely to outweigh uncertain benefits, significantly hindering research and development efforts. The existence of applications where broad consensus exists that simulating opinions is desirable, such as when human survey subjects would be exposed to harmful content [17] or trade secrets [27], motivates the need for approaches that lower the cost of early-stage exploration of LLM usage to simulate opinions.
Our proposal: early-stage, low-cost quality checks
We introduce three diagnostic tests that probe desirable structural properties of simulated opinions:
@@
    Prompt Stability Across Model Updates—whether outputs of prompts remain relatively consistent over time across minor model updates;
@@
These tests can be conducted prior to any large-scale human validation efforts, which would only be warranted in cases where an LLM simulation approach yields positive results on all three diagnostic tests.
Validating opinions simulated with LLMs shares challenges with validating subjective human annotations. When humans supply those opinions, researchers rely on long-standing quality controls to decide whether the raw or aggregated annotations are trustworthy or reliable [28, 29, 30], and whether annotators were paying attention and comprehending the survey questions [31]. Thus, our quality checks are informed by research on assessing quality standards and biases in human subjective data annotation [31, 30, 32, 28, 33, 34, 35, 36, 29]. Our proposed tests build on this rich literature, extending it to the evaluation of LLM-simulated opinions with a particular focus on Likert-scale opinion data. The first two tests rely solely on model outputs, while the third can be instantiated with or without human data.
@@
We show that, despite their simplicity, these proposed quality control tests reveal several failure modes of the common backstory prompting approach used with several state-of-the-art LLMs. Thus, these tests offer practitioners a triage stage that filters out weak prompts or models before committing serious resources. We conclude by discussing implications of the observed failure modes and recommend how organizations can utilize and embed our proposed tests in their risk management practices when considering the use of LLMs for opinion simulation. We make our crowdsourced dataset of claims with human and LLM annotations (TopicMisinfo) publicly available for future research.
Crafting a domain-specific testbed for quality checks
To demonstrate the potential of our proposed quality checks, we construct a domain-specific testbed focused on content moderation, specifically the prioritization of claims—i.e., potential harmful misinformation—for human review by fact-checkers. In determining whether something should be prioritized for fact-checking, factors such as the likelihood that a claim is harmful to a group if false (e.g. claims that can foster anti-immigrant sentiments) are important considerations [37, 38]. This prioritization process, traditionally carried out by the fact-checkers themselves, is inherently opinion-driven and resource-intensive [39], making it an ideal scenario for exploring the utility of simulating opinions.
A testbed designed to evaluate group-level opinion simulation should include both cases where the groups are expected to agree and cases where they are expected to be sharply divided. Among demographic dimensions that offer strong prior expectations, gender opinions are particularly well-documented [40, 41]. We focus particularly on the prioritization criterion of “potential harm to specific groups”, as prior work has indicated that gender-level differences will be particularly salient along this axis of consideration [40, 41].
We built TopicMisinfo, a 160-claim dataset spanning topics that can be expected to elicit gender disagreement and consensus. The selection of these topics is anchored in findings from the American National Election Studies (ANES) [42], which helps us craft a diversity in topics that are likely to provoke varying degrees of gendered disagreement. We also included “gold” comprehension checks, which are claims that are harmless and either obviously true (e.g., “A circle is round.”) or obviously false (e.g., “The tallest tree on Earth touches Mars.”), with a very strong prior on gender-level consensus of opinion.
We apply our diagnostic tests to simulated outputs from several commercially available LLMs: GPT-3.5-Turbo (checkpoints 06-13-23, 11-06-23, and 01-25-24), GPT-4, LLaMA 3, Titan-Text-Premier, and Mistral-Large, referred to hereafter as GPT3.5a, GPT3.5b, GPT3.5c, GPT-4, LLaMA3, Titan, and Mistral, respectively. Following established gender-conditioning methods [15, 22], we elicit group-level responses through “backstory” prompts that instruct models to adopt personas of men or women. We also collect a small sample of task-specific human opinion data and show how small amounts of human annotations–if available–can be incorporated into our quality checks, which we do in Experiment 3b. Detailed descriptions of dataset composition, example claims, prompting templates, and descriptive human annotation statistics are provided in Supplementary Materials S1–S3, and we make the TopicMisinfo dataset publicly available to facilitate further research [^1] It will be made available for camera-ready version in an online repository..
@@ 2 Results
In this section, we further explain the three proposed tests, and analyze the results of deploying them for evaluating LLM performance on a particular opinion task related to prioritization of claims for fact-checking: assessing the potential harm of claims to certain groups as a means to prioritize it for review [^2] While not the only criteria relevant for prioritization, research has noted that this is an important axis of consideration used by fact-checkers [38, 39].. These tests, designed to emphasize logical consistency, prompt stability, and alignment with stakeholder expectations, serve as quality checks that provide crucial information in early-stage deliberation when considering the use of LLMs to simulate group-level opinions and to develop more reliable prompting techniques.
@@
Figure 1: Experiment 1 Results. Plausible ranges of $q_{0}$ across topics for different models are highlighted in blue. Bonferroni correction applied to account for high likelihood of type I error due to multiple hypothesis testing ($\alpha^{*}=0.05/20=0.0025$).
Testing for logical consistency can be an effective way to gauge the reliability of human annotators [35] or data aggregation methods [30]. For instance, Arguete et al. [35] use logical contradiction as a measure of human annotator inattention. We introduce a related methodology for evaluating logical consistency of LLMs in group-level opinion simulation by testing whether a model, when prompted for an “average" or “neutral” perspective on a Likert-scale assessment, generates a stance that is a convex combination of the synthetic group-level opinions that it simulates. If a model’s average opinion response is not positioned plausibly within the convex hull of divergent group opinions, it highlights a statistical impossibility that suggests a critical logical inconsistency in the underlying approach to generating viewpoints.
In this paper, we define statistical neutrality as an opinion located within the convex hull of all group-level opinions within a social category (e.g., gender, race, etc.), including the extrema (i.e., only representing one group). This is distinct from neutrality focused on disinterest or impartiality [43, 44]. In the absence of any single established approach to prompting an LLM to generate a “statistically neutral” viewpoint, four distinct prompts were evaluated, each with different approaches to querying statistical neutrality. See Supplementary Materials S3 for exact prompt specifications.
@@
Note that this test for logical consistency is fundamentally different from assessments of consistent reasoning discussed in prior work [45, 46] in that the current test focuses on whether the positioning of neutral or average opinions is statistically reasonable. Statistical plausibility is essential because, in the absence of such consistency, one of those simulated sets of opinions must be wrong.
@@
Experiment 2: Prompt Stability Over Time
Measuring the stability of a single LLM’s outputs for the same prompt across minor model updates (or “checkpoints”) offers critical insight into its reliability. This kind of test is motivated by quality control in human annotation pipelines, where inconsistent judgments cast doubt on an annotator’s credibility [34]. As Abercrombie et al. [32, pg.1] observe, “consistency and label stability are important because, without them, annotation schemes are unlikely to be repeatable or reproducible.”
Prompt stability is also closely tied to the concept of backwards compatibility, a long-standing concern in software engineering [47], and one of growing importance when we consider the use of “off-the-shelf” machine learning models applied to a variety of application domains [48]. In AI systems, backwards compatibility refers to whether updated models make categorically similar predictions or errors to previous versions. For LLMs, it is reasonable to expect backwards compatibility within a given “model class” (e.g., across checkpoints of GPT-3.5-Turbo), though not necessarily between model classes (e.g., GPT-3.5-Turbo vs. GPT-4). Updates within a model class are generally intended to improve safety and reliability, not to introduce significant behavioral or quality changes [49]. Yet, emerging evidence suggests this expectation is not always met. For example, Chen et al. [50] found that two checkpoints of GPT-3.5-Turbo disagreed on 27% of multiple-choice responses in the OpinionQA dataset.
@@
Figure 2: Experiment 2 Results. This chart highlights statistically significant changes in model outputs across three updates (June ‘23, Nov ‘23, and Jan ‘24) for three types of prompts: neutral, men-conditioned, and women-conditioned. Stars mark statistical significance (* p$<$0.05; **∗ ∗ p$<$0.01; ***∗ ∗ ∗ p$<$0.001).
@@
Ratings of these quality control items should remain stable and low over time. All of these claims were designed to be harmless truths or lies. However, we see small but significant differences in ratings across checkpoints for men-conditioned ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 0.19, p $<$ 0.001), neutral ($\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5c) = 0.55, p $<$ 0.001), and women-conditioned prompts ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 0.51, p $<$ 0.001).
@@
Topic: Gold Topic: Abortion
(Prior: Insig. Differences) (Prior: Sig. Differences)
GPT-3.5a 0.25*** 2.09***
GPT-3.5b 0.20* 1.09***
GPT-3.5c 0.38** 1.27***
GPT-4 0.00 0.07
llama3 0.13 0.06
mistral 0.02 0.26
titan 0.19** 0.30*
Table 1: Experiment 3a — Alignment between model-predicted gender differences and stakeholder priors. Each entry is the absolute mean difference in predicted gender-level responses. Stars mark statistical significance (* p$<$0.05; **∗ ∗ p$<$0.01; ***∗ ∗ ∗ p$<$0.001). Blue values agree with the prior; red values contradict it. "Gold" claims are those that are harmless and unambiguously true of false (e.g., "A circle is round") that would be used as a comprehension check for human annotators. Results indicate that no model simultaneously satisfies both priors.
@@
Stakeholders that rely on LLM-simulated opinions for decision making have a vested interest in verifying that LLMs accurately reflect group-level viewpoints. If an LLM does not position group-level opinions appropriately, it risks introducing biases into human decision-making. For example, if fact-checking systems rely on LLMs to estimate which claims different groups are likely to find harmful, a model that underrepresents gender differences in concern may deprioritize fact-checks on topics that disproportionately affect one group’s information environment. Notably, even in the absence of “ground truth" human data to validate a particular opinion survey task, stakeholders often have valuable domain knowledge that can be leveraged to validate the reliability of LLMs. In this experiment, we focus on the positioning of gender differences in opinion, asking: Does our approach to opinion simulation position men’s and women’s opinions in ways that align with stakeholder expectations regarding the degree of disagreement across different topics?
We define “expectations” by leveraging topics where prior research, survey results, or common sense strongly indicates minimal or substantial disagreement between gender groups. Experiment 3 examines two complementary scenarios: (a) whether the model violates well-established priors from common sense, domain knowledge, or survey data—e.g., that a neutral statement like “A circle is round” should not elicit divergent opinions by gender, whereas a claim about abortion might, drawing inspiration from prior comprehension-based quality control checks [33]; and (b) whether the model distorts stakeholder expectations on less obviously gender-polarized issues (e.g., politics, immigration, health and science), by introducing task-specific, small sample human survey data that serves as a reasonable prior for stakeholder expectations. This two-part design allows us to identify when a model introduces artificial rifts or masks genuine ones, both crucial factors in gauging an LLM’s trustworthiness for representing group-specific viewpoints. It also allows us to illustrate how the third quality check may be adapted depending on the type of prior information available. See Materials and Methods for details on the methodology for Experiments 3a and 3b. Tables 1 and 2 highlight the results for Experiments 3a and 3b respectively, and we provide a replication experiment of Experiment 3b with a prompt variant in Supplementary Materials S6.
@@
To test for exaggeration of opinion differences, we compared the output of LLMs prompted to reflect the opinions of men and women on topics where we had very strong priors that no division should exist. For instance, the Gold topic in the TopicMisinfo dataset contains a set of items that are harmless and uncontroversially true (e.g., “A circle is round”) or false (e.g., “The tallest tree on Earth creates a land bridge to Mars”), and they were compiled and used as a means to make sure that human annotators were paying attention and comprehending the task. We would very reasonably expect no opinion disagreement on claims of this type. However, we see GPT3.5a ($\hat{L}_{GPT3.5a}=0.25$, p $<$ 0.001), GPT3.5b ($\hat{L}_{GPT3.5b}$ = 0.20, p $<$ 0.05), and GPT3.5c ($\hat{L}_{GPT3.5c}$ = 0.38, p $<$ 0.01), as well as Titan ($\hat{L}_{Titan}$ = 0.19, p $<$ 0.01), tended to produce statistically significant differences in opinions by gender for claims on this topic. See Supplementary Materials S5 for specific examples of exaggeration.
@@
When preliminary data from real human responses are available, this can inform the stakeholder prior. For this purpose, we use crowdsourced annotations of the TopicMisinfo dataset. When comparing LLM responses to crowdsourced human responses, GPT3.5a, GPT3.5b, and GPT3.5c tended to exaggerate differences between groups, particularly for topics that are not typically considered divisive between genders. For example, claims from topics such as entertainment ($\hat{L}_{GPT3.5a}=0.84$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.49$, $p<0.01$; $\hat{L}_{GPT3.5c}=0.55$, $p<0.001$), health & science ($\hat{L}_{GPT3.5a}=1.05$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.52$, $p<0.01$), U.S. politics ($\hat{L}_{GPT3.5a}=1.09$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.58$, $p<0.001$; $\hat{L}_{GPT3.5c}=0.36$, $p<0.01$), and weather & climate ($\hat{L}_{GPT3.5a}=0.65$, $p<0.01$; $\hat{L}_{GPT3.5b}=0.62$, $p<0.01$; $\hat{L}_{GPT3.5c}=0.46$, $p<0.05$) saw significant overestimation of group-level disagreement by these models.
@@
To test for erosion of opinion differences, we compared the output of LLMs prompted to reflect the opinions of men and women on topics where we had very strong priors that division should exist. For instance, it is well-established that men and women may perceive harms differently by claims about abortion and women’s health [41, 40], and this topic is included in the TopicMisinfo dataset. Thus, we have some prior that a reasonable LLM should display differences in outputs when conditioned to respond as if it were either a man or woman. However, we didn’t find any significant differences between the outputs from men- or women-conditioned prompts on the topic for Mistral ($\hat{L}_{Mistral}=0.26$, p = 0.09), GPT-4 ($\hat{L}_{GPT4}=0.06$, p = 0.16), and LLaMa3 ($\hat{L}_{LLaMa}=0.06$, p = 0.23). See Supplementary Materials S5 for specific examples of erosion.